OpenStudy (anonymous):

(x+0,2)^2-(x-0,2)^2=? The book says the answer is 0.8x but I cant seem to make it work on paper... Please help

OpenStudy (anonymous):

\[ x^{2}+0.04-x^2+0.04\].........0.04+0.04=0.08....thats all i know

OpenStudy (anonymous):

x^2 + 0.4x + 0.04 - ( x^2 - 0.4x + 0.04 ) = ? x^2 - x^2 + 0.4x + 0.4x +0.04 -0.04 = 0.8x

OpenStudy (anonymous):

Thank you, it looks good to me.